Find the value of k if the straight line 6x-10y+3=0,kx-5y+8=0 are parallel

Grade:11

1 Answers

Debarati
2 years ago
Parallel lines means that this system of equation has no solution.
So, a1/a2=b1/b2\neqc1/c2
=> 6/k=-10/-5\neq3/8
=> 6/k= 2
=> k= 3
\thereforek= 3

I hope you understood.